OpenStudy (anonymous):

Rectangular box has volume x 3 − 8 cubic inches. If the height of the box is x − 2 inches, what is the area of the base of the box, in square inches? V=b*h

OpenStudy (anonymous):

rectangular box has volume x^3 − 8 cubic inches. If the height of the box is x − 2 inches, what is the area of the base of the box, in square inches?

Mehek (mehek14):

\(x^3-8=b*(x-2)\) you are going to solve for b \(\dfrac{x^3-8}{x-2}=b\) \(\dfrac{x^3}{x}=x^2 \\ \dfrac{-8}{-2}=4 \\ x^2+4 = b\)

Mehek (mehek14):

Did you understand or no? @TMA76